### v1

@@ Abstract
Context: Deep learning has achieved remarkable progress in various domains. However, like any software system, deep learning systems contain bugs, some of which can have severe impacts, as evidenced by crashes involving autonomous vehicles. Despite substantial advancements in deep learning techniques, little research has focused on reproducing deep learning bugs, which hinders resolving them. Existing literature suggests that only 3% of deep learning bugs are reproducible, underscoring the need for further research.
Objective: This paper examines the reproducibility of deep learning bugs. We identify edit actions and useful information that could improve deep learning bug reproducibility.
Method: First, we construct a dataset of 668 deep learning bugs from Stack Overflow and Defects4ML across three frameworks and 22 architectures. Second, out of the 668 bugs, we select 102 bugs using stratified sampling and try to determine their reproducibility. While reproducing these bugs, we identify edit actions and useful information necessary for their reproduction. Third, we used the Apriori algorithm to identify useful information and edit actions required to reproduce specific bug types. Finally, we conduct a user study with 22 developers to assess the effectiveness of our findings in real-life settings.
Results: We successfully reproduced 85 out of 102 bugs attempted. We identified ten edit actions and five useful information categories that can help us reproduce the deep learning bugs. Our findings improved bug reproducibility by 22.92% and reduced reproduction time by 24.35% based on a user study with 22 developers.
@@ 1 Introduction
Like any software system, deep learning systems are prone to bugs. Bugs in deep learning systems could arise from various sources, including errors in the dataset, incorrect hyperparameters, and incorrect structure of the deep learning model BRAIEK2020110542 . These bugs could lead to program failures, poor performance, or incorrect functionality, as reported by existing literature islamfse19 . They can also lead to serious consequences, as shown by the fatal crash involving Uber’s self-driving car selfdrivingcarcrash . Thus, we must fix the bugs before deploying a deep learning model in production. However, one must reproduce the bugs before fixing them to verify their presence/absence in the system. Unfortunately, reproducing DL bugs is challenging due to deep learning systems’ multifaceted nature and dependencies, which encompass data, hardware, libraries, frameworks, and client programs. Furthermore, deep learning systems are inherently non-deterministic (i.e., random weight initialization), which leads to different outcomes across multiple runs and thus makes the reproduction of deep learning bugs challenging nagarajan2018impact . Moreover, they also suffer from a lack of interpretability krishnan2020against , which makes the reproduction of deep learning bugs challenging in comparison to traditional software bugs. According to existing investigations defects4ml , only 3% of DL bugs are reproducible, further demonstrating the challenges in reproducing DL bugs.
@@
In this paper, we conduct an empirical study to better understand the challenges in reproducing deep learning bugs. First, we collect a total of 568 DL bugs from Stack Overflow posts. Then, we extend them with 100 DL bugs from the benchmark dataset of Moravati et al. defects4ml , which makes up our final dataset of 668 DL bugs. Using the taxonomy of Humbatova et al. defects4ml , we divide these bugs into five categories: 167 model, 213 tensor, 145 training, 113 GPU, and 30 API bugs. Second, by using stratified sampling, we select 102 of these bugs and determine the reproducibility status of these bugs by attempting to reproduce them using the code snippet and complementary information from Stack Overflow and the benchmark dataset. Third, the first author manually analyzed the produced artefacts to determine the categories of information that are useful for deep learning bug reproduction. Fourth, we use the information gathered during the bug reproduction and create a dataset of transactions that link the type of bug with edit actions and information categories. Then, our study establishes connections among the information categories, edit actions, and the type of bugs by employing the Apriori algorithm apriori . Finally, to further validate our findings, we conducted a developer study with 22 participants from industry and academia. Half of the participants were asked to reproduce bugs using our identified information, while the other half were asked to reproduce the same bugs without access to that information. Results from the user study show that our recommended edit actions and information categories can reduce the time to reproduce the bug by 24.35%. Thus, we answer three important research questions as follows:
@@
Determining the key edit actions that are crucial for reproducing deep learning bugs is important since almost none of the bugs can be reproduced using the verbatim code snippet. By manually reproducing 85 deep learning bugs, we identify ten key edit actions that could be useful to reproduce deep learning bugs (e.g., input data generation, neural network construction, hyperparameter initialization). -
RQ2: What information categories and edit actions are useful for reproducing specific types of deep learning bugs?
@@
RQ3: Do the suggested edit actions and identified information categories contribute to enhancing the reproducibility of deep learning bugs in practice?
To assess the effectiveness of our suggested edit actions and identified information categories in improving bug reproducibility, we conducted a user study involving 22 professional developers. In our developer study, the participants assigned to the experimental group used our suggested edit actions and information categories to reproduce deep learning bugs. In contrast, the participants in the control group reproduced the bugs without access to the suggested edit actions and information categories. Based on the feedback from the developers, we found that our recommended edit actions and highlighted information categories (a) helped them reproduce 22.92% more bugs and (b) decreased their time to reproduce the deep learning bugs by 24.35%.
The remainder of the paper is organized as follows. Section 2 presents a motivating example that highlights the challenges in reproducing deep learning bugs using existing techniques. Section 3 then describes our study methodology in detail, including data collection, dataset construction, environment setup, qualitative analysis, and user study design. Next, Section 4 presents the findings by answering our three research questions. It summarizes the key edit actions, useful information categories, and their relationships with bug types that enhance deep learning bug reproducibility. Section 5 and Section 6 provide a discussion of our results, including a proof-of-concept tool, comparison with language models and threats to validity. Finally, Section 7 reviews related literature, and Section 8 concludes the paper.
@@ 2 Motivating Example
Bug reports or programming Q&A posts might not always provide sufficient information to reproduce a deep learning bug. Let us consider the example question shown in Fig. LABEL:fig:irreproducibleMotivatingExample motivatingExample1 . Here, the user attempts to train a neural network using Keras. The input data type is a object, and the user aims to pass the training dataset to the method. However, s/he discovers that their object is not recognized by the Keras library. The user also provides a code snippet to aid the reproducibility of the bug. Unfortunately, the issue cannot be reproduced since the provided information does not contain any dataset for training, network definition, hyperparameters, or required dependencies. In Stack Overflow, this question has failed to receive a precise response. Even though the above code (Fig. LABEL:fig:irreproducibleMotivatingExample) could be made compilable or executable using edit actions, the bug cannot be reproduced due to its complex nature. Since the earlier study msr2019mondal does not deal with any deep learning bugs, their suggested edits might also not be effective.
@@
Let us consider another example question shown in Fig. LABEL:fig:motivatingExample motivatingExample2 . Here, the user wants to obtain the confusion matrix from a multi-class classification model. Unfortunately, s/he runs into a runtime error, as the method does not support multiple output labels. To reproduce the bug, we first determine the type of this bug. By manually analyzing the Stack Overflow post, code snippet, and the corresponding stack trace, we found that the error was related to the incorrect shape of data being passed to the method. Since the bug was triggered by a wrong shape of the training dataset, it falls into the ‘Tensor’ category according to the taxonomy of Humbatova et al. taxonomyRealFaults . Then, we check the useful information suggested by our study (see RQ2) for reproducing tensor bugs. The top 3 items are Data, Logs, and Code Snippet, as highlighted in Table. 7. Although the question in Fig. LABEL:fig:motivatingExample contains the relevant code snippet and logs, it does not contain any information about the dataset for training, which is the most useful information for reproducing tensor bugs according to our findings (RQ2). Subsequently, we also check the recommended edit actions to reproduce a tensor bug. From Table 8, we use the top 3 edit actions – Input Data Generation, Import Addition, and Hyperparameter Initialization. When these edit actions and information categories were applied to the code, we were able to successfully reproduce the bug, which indicates the effectiveness of our suggested actions and information. Similarly, this issue was reproduced by other users of Stack Overflow, and the question received a correct solution within two hours of its submission, as shown in Fig. LABEL:fig:AcceptedAnswer.
@@ 3.1 Selection of Data Sources
We select Stack Overflow as a primary data source for our study. It is the largest programming Q&A site for programming topics containing over 24 million questions and 35 million answers SOUsageData . Thus, Stack Overflow could be a potential source for deep learning (DL) bugs. Developers often submit their encountered problems on Stack Overflow, when building their deep learning applications. According to a recent work zhao2021state , Stack Overflow contains at least 30 topics and 80K posts related to deep learning issues, which makes it a potential source for our data. We also select the benchmark dataset of Moravati et al. defects4ml containing 100 DL bugs as our second source of data.
@@ 3.2 Dataset Construction
The above filtration step resulted in 3,412 Stack Overflow posts. To construct our final dataset, we apply another filter to these posts where we leverage the tags from Humbatova et al. taxonomyRealFaults . Our idea was to select representative samples from different types of deep learning bugs. To reduce the risk of miscategorization, we thus only selected the tags in the taxonomies and sub-taxonomies proposed by Humbatova et al. Table. 1 shows our selected 3 tags for model bugs, 1 for tensor bugs, 11 for training bugs, 3 for GPU bugs, and 7 for API bugs. We look for these tags in the above Stack Overflow posts and collect 113 Model bugs, 193 Tensor Bugs, 95 Training Bugs, 101 GPU Bugs, and 24 API Bugs. We also discover 42 bugs belonging to multiple categories (e.g., 5 Model & Tensor Bugs, 5 Tensor & GPU Bugs, 3 Training & API Bugs). Thus, the final dataset contains a total of 668 bugs (568 from our dataset + 100 from Defects4ML defects4ml ) and captures a balanced representation from different types of DL bugs, as shown in Table 2 (Step 3, Fig. 1).
@@ 3.4 Qualitative Analysis
We follow a two-step approach to reproduce the deep learning issues reported in the studied Stack Overflow posts. First, we gather complementary information about the issue, such as the dataset, library versions, and the framework used. Second, we attempt to reproduce the issue using the code snippet and supporting data (e.g., dataset information, environment configurations, hyperparameters, and training logs). Following the successful reproduction of a bug, we capture various information related to each bug, such as the deep learning architecture involved, edit actions used to reproduce the issue, time taken to reproduce the issue, type of issue reproduced, and the type of information present in the Stack Overflow post. All these data gathered during issue reproduction will help us determine the useful information and edit actions needed to reproduce specific types of deep learning bugs (Step 4, Fig. 1).
| Edit Action | Overview |
A: Input Data Generation (D) |
Generating input data which simulates the data used for training the model. |
A: Neural Network Construction (N) |
Reconstructing or modifying the neural network based on the information provided. |
A: Hyperparameter Initialization (H) |
Initializing the hyperparameters for training, such as batch size and number of epochs. |
A: Import Addition and Dependency Resolution (R) |
Determining the dependencies in the code snippet and adding the missing import statements. |
A: Logging (L) |
Adding appropriate logging statements to capture relevant information during reproduction. |
A: Obsolete Parameter Removal (O) |
Removing outdated parameters or functions to match the parameters of the latest library versions. |
A: Compiler Error Resolution (C) |
Debugging and resolving compiler errors that arise due to syntactic errors in the provided code snippet. |
A: Dataset Procurement (P) |
Acquiring the necessary datasets and using them to train the model. |
A: Downloading Models & Tokenizers (M) |
Fetching pre-trained models and tokenizers from external sources. |
A: Version Migration (V) |
Updating code to adapt to changes introduced in newer library or framework versions. |

### v2

@@ Abstract
Context: Deep learning has achieved remarkable progress in various domains. However, like any software system, deep learning systems contain bugs, some of which can have severe impacts, as evidenced by crashes involving autonomous vehicles. Despite substantial advancements in deep learning techniques, little research has focused on reproducing deep learning bugs, which is an essential step for their resolution. Existing literature suggests that only 3% of deep learning bugs are reproducible, underscoring the need for further research.
Objective: This paper examines the reproducibility of deep learning bugs. We identify edit actions and useful information that could improve the reproducibility of deep learning bugs.
Method: First, we construct a dataset of 668 deep learning bugs from Stack Overflow and GitHub across three frameworks and 22 architectures. Second, out of the 668 bugs, we select 165 bugs using stratified sampling and attempt to determine their reproducibility. While reproducing these bugs, we identify edit actions and useful information for their reproduction. Third, we used the Apriori algorithm to identify useful information and edit actions required to reproduce specific types of bugs. Finally, we conduct a user study involving 22 developers to assess the effectiveness of our findings in real-life settings.
Results: We successfully reproduced 148 out of 165 bugs attempted. We identified ten edit actions and five useful types of component information that can help us reproduce the deep learning bugs. With the help of our findings, the developers were able to reproduce 22.92% more bugs and reduce their reproduction time by 24.35%.
@@ 1 Introduction
Like any software system, deep learning systems are prone to bugs. Bugs in deep learning systems could arise from various sources, including errors in the dataset, incorrect hyperparameters, and incorrect structure of the deep learning model BRAIEK2020110542 . These bugs could lead to program failures, poor performance, or incorrect functionality, as reported by existing literature islamfse19 . They can also lead to serious consequences, as shown by the fatal crash involving Uber’s self-driving car selfdrivingcarcrash . Thus, we must fix the bugs before deploying a deep learning model in production. However, one must reproduce the bugs before fixing them to verify their presence/absence in the system. Unfortunately, reproducing DL bugs is challenging due to deep learning systems’ multifaceted nature and dependencies, which encompass data, hardware, libraries, frameworks, and client programs. Furthermore, deep learning systems are inherently non-deterministic (i.e., random weight initialization), which leads to different outcomes across multiple runs and thus makes the reproduction of deep learning bugs challenging nagarajan2018impact . Moreover, they also suffer from a lack of interpretability krishnan2020against , which makes the reproduction of deep learning bugs challenging in comparison to traditional software bugs. According to existing investigations defects4ml , only 3% of their analyzed deep learning bugs were reproducible, further demonstrating the challenges in reproducing DL bugs.
@@
In this paper, we conduct an empirical study to better understand the challenges in reproducing deep learning bugs. First, we collect a total of 568 DL bugs from Stack Overflow posts. Then, we extend them with 100 DL bugs from the benchmark dataset of Moravati et al. defects4ml , which makes up our final dataset of 668 DL bugs. Using the taxonomy of Humbatova et al. defects4ml , we divide these bugs into five categories: 167 model, 213 tensor, 145 training, 113 GPU, and 30 API bugs. Second, by using stratified sampling, we select 165 of these bugs and determine their reproducibility status by attempting to reproduce them using the code snippet and complementary information from Stack Overflow and the benchmark dataset. Third, the first author manually analyzed the produced artefacts to determine the categories of information that are useful for deep learning bug reproduction. Fourth, we use the information gathered during the bug reproduction and create a dataset of transactions that link the type of bug with edit actions and component information. Then, our study establishes connections among the component information, edit actions, and the type of bugs by employing the Apriori algorithm apriori . Finally, to further validate our findings, we conducted a developer study with 22 participants from industry and academia. Half of the participants were asked to reproduce bugs using our identified information, while the other half were asked to reproduce the same bugs without access to that information. Results from the user study show that our recommended edit actions and component information can reduce the time to reproduce the bug by 24.35%. Thus, we answer three important research questions as follows:
@@
Determining the key edit actions that are crucial for reproducing deep learning bugs is important since almost none of the bugs can be reproduced using the verbatim code snippet. By manually reproducing 148 deep learning bugs, we identify ten key edit actions that could be useful to reproduce deep learning bugs (e.g., input data generation, neural network construction, hyperparameter initialization). -
RQ2: What types of component information and edit actions are useful for reproducing specific types of deep learning bugs?
@@
RQ3: How do the suggested edit actions and information affect the reproducibility of deep learning bugs?
To assess the effectiveness of our suggested edit actions and component information in improving bug reproducibility, we conducted a user study involving 22 professional developers. In our developer study, the participants assigned to the experimental group used our suggested edit actions and component information to reproduce deep learning bugs. In contrast, the participants in the control group reproduced the bugs without access to the suggested edit actions and component information. We found that our recommended edit actions and component information (a) helped the developers reproduce 22.92% more bugs and (b) decreased their time to reproduce the deep learning bugs by 24.35%.
The remainder of the paper is organized as follows. Section 2 presents a motivating example that highlights the challenges in reproducing deep learning bugs. Section 3 introduces the necessary background information for deep learning bugs and their bug reports. Section 4 then describes our study methodology in detail, including data collection, dataset construction, environment setup, qualitative analysis, and user study design. Next, Section 5 presents the findings by answering our three research questions. It summarizes the key edit actions, component information, and their relationships with bug types that enhance deep learning bug reproducibility. Section 6 discusses our results and shows how our findings can be used to improve LLMs. Section 7 discusses the threats to the validity of this study. Finally, Section 8 reviews related literature, and Section 9 concludes the paper.
@@ 2 Motivating Example
Bug reports or programming Q&A posts might not always provide sufficient information to reproduce a deep learning bug. Let us consider the example question shown in Fig. LABEL:fig:irreproducibleMotivatingExample motivatingExample1 . Here, the user attempts to train a neural network using Keras. The input data type is a object, which is used as the base object for fitting a sequence of data, such as a dataset sequence The user aims to pass the training dataset as a object to the method. However, s/he discovers that their object is not recognized by the Keras library. The user also provides a code snippet to aid the reproducibility of the bug. Unfortunately, the issue cannot be reproduced since the provided information does not contain the required dependencies and imports. In Stack Overflow, this question has failed to receive a precise response. Even though the above code (Fig. LABEL:fig:irreproducibleMotivatingExample) could be made compilable or executable using edit actions, the bug cannot be reproduced due to its complex nature. Since the earlier study msr2019mondal does not deal with any deep learning bugs, their suggested edits might also not be effective.
@@
Let us consider another example question shown in Fig. LABEL:fig:motivatingExample motivatingExample2 . Here, the user wants to obtain the confusion matrix from a multi-class classification model. Unfortunately, s/he runs into a runtime error, as the method does not support multiple output labels. To reproduce this bug, we first generate a synthetic multi-class dataset since the training data is missing from the question. We also add the required import statement for the ‘confusion_matrix’ function. Then, we initialize the model hyperparameters based on the code snippet. When we run the code on this synthetic data, it triggers the same runtime error due to the wrong data shape being passed to ‘confusion_matrix()’. Thus, by extracting the key information and applying the necessary edit actions, we were able to reproduce the bug. Similarly, this issue was reproduced by other users of Stack Overflow, and the question received a correct solution within two hours of its submission, as shown in Fig. LABEL:fig:AcceptedAnswer. To summarize, by generating a synthetic dataset, adding the necessary import statement, and initializing the model hyperparameters, we can reproduce the example deep learning bug. Our study in this article proposes the methodology to extract the information and determine the edit actions necessary for systematically reproducing deep learning bugs.
3 Background
3.1 Deep Learning Bugs
Developers often encounter different types of bugs when they write their deep-learning software. According to existing literature taxonomyRealFaults ; islam2019fse , these bugs can be classified into five categories: Training, Model, Tensor and Input, API, and GPU. Table. 1 shows the prevalence of these bugs in deep learning systems. We discuss these bugs in detail below.
Training Bugs: Training bugs encompass a wide range of issues that impact various aspects of the training process, including data preprocessing, hyperparameter tuning, loss function selection, and model optimization. Data preprocessing steps, such as normalization and input scaling, are frequently overlooked or improperly executed, leading to subpar results. Quality of training data is another critical factor that can affect of the training process croft2023data . Insufficient, imbalanced, or incorrectly labelled training data can significantly impact the model’s ability to generalize and produce accurate predictions. Hyperparameter problems are quite common, with issues such as incorrect learning rates, inappropriate batch sizes, and inadequate selection of the number of epochs taxonomyRealFaults . Similarly, the selection and implementation of loss functions could be suboptimal or flawed. The choice and tuning of optimizers can also pose challenges, as incorrect selections or settings can hinder the model performance. Additional problems that may arise during training include memory mismanagement and the absence of data augmentation techniques. Overall, the multifaceted nature of neural network training, with its numerous hyperparameters and intricate architectures, often introduces bugs into the training process.
Model Bugs: Model bugs stem from the neural network architectures and their components. Common faults include wrong model type for a task, incorrect network depth with too few or too many layers, and issues with specific layers. For instance, choosing a multilayer perceptron over a convolutional neural network for an image classification task leads to poor performance, as the model type does not fit the problem. Similarly, networks with inadequate or excessive depth may fail to capture patterns or may overfit, respectively. Faulty layer design, such as the omission of data manipulation layers, could lead to flaws in feature extraction, whereas redundant layers increase the complexity of the model. Other frequent sources of bugs are suboptimal layer parameters like neuron counts, kernel dimensions, and stride lengths. Finally, inaccurate activation function selection, such as incorrectly using sigmoid instead of ReLU in the hidden layers, can degrade the model performance. Overall, model bugs stem from improper network design and configuration, leading to a fundamentally unsuitable model structure.
Table 1: Prevalence of Different Types of Deep Learning Bugs Type of Bug Training Model Tensor and Input API GPU Prevalence 52.5% 19.7% 19.5% 5.3% 2.9% -
Tensor & Input Bugs: Tensor bugs stem from incorrect tensor shapes provided in the deep learning models, i.e. mismatches between the expected tensor dimensions and the actual shape of the input tensor. Potential causes include errors in padding, indexing, or tensor transformations like transposing before model ingestion. Input bugs occur when the input data does not match the specifications the model requires. Examples include incorrect data types (e.g. strings instead of numeric), inconsistent data shapes (e.g. wrong matrix dimensions), and improper data formats (e.g. channels reversed in images). While tensor bugs arise from incorrect shaping of the input data, input bugs stem from invalid or incompatible characteristics of the raw input itself. Both categories can seriously hurt the model performance by propagating flawed representations, even if no crashes occur.
API Bugs: API bugs primarily occur due to an incorrect usage of the API. Several factors can lead to incorrect API usage. First, API definitions often change across framework versions, causing developers to use outdated or incompatible syntax. Second, there can be API incompatibility within the same framework. For example, in TensorFlow 1.x, the developers could use the Estimator API with model_fn() to build models. But, in TensorFlow 2.0, Estimators were removed, so code written for v1 would cause compile-time errors when the framework was updated. Finally, unclear or confusing documentation can result in developers misunderstanding the proper use of an API. This can lead to multiple issues, such as using an API incorrectly in a manner that deviates from the framework’s intended use. Furthermore, developers may encounter challenges when the documentation lacks clear specifications of crucial API calls for a workflow or misplaces API call explanations within the code.
GPU Bugs: GPU bugs are the most difficult bugs to localize and reproduce, as observed by Jahan et al. jahan2024towards . GPU bugs can occur due to various factors such as incorrect GPU device referencing, failed parallelism, improper state sharing between subprocesses, and faulty data transfers to the GPU. Incorrect device referencing stems from developers misunderstanding the mapping between CPU and GPU environments sanders2010cuda . Failed parallelism and improper state sharing often result from subtle data races or synchronisation errors, respectively aviram2012efficient ; zhang2018g . Faulty data transfers arise from mismatches between CPU and GPU data formats tiwari2015understanding . Overall, the domain-specific nature of GPU programming, the rapid evolution of GPU architectures, and the lack of sufficient validation could lead to difficult-to-diagnose GPU bugs.
3.2 Deep Learning Bug Reports
To date, only limited investigation has been conducted to understand the patterns of deep learning (DL) bug reports. Long et al. long2022reporting performed the first exploratory study to analyze the bug reporting trends and patterns for deep learning frameworks. Their work revealed several key themes that provide insight into the nature and lifecycle of these bug reports:
Root Causes and Prevalence: Their study has found that low training speed is the most common symptom for submitting performance-related bug reports, ranging from 27% to 67% across the different DL frameworks. However, no consistent pattern was observed for the root causes of accuracy-related reports. This suggests that performance issues, especially those manifesting as slow execution speed, are a major pain point encountered by the users of DL frameworks.
Affected Stages of DL Pipeline: Across the frameworks studied, the training stage was found to be the most prevalent in performance and accuracy bug reports, ranging from 38% to 77%. This finding is not surprising, as training is typically the most computationally intensive and time-consuming stage of the DL pipeline shi2016benchmarking . Performance bottlenecks or accuracy issues encountered during training can significantly impact the overall usability and efficiency of the framework.
Report Quality and Resolution: Their study found that a majority of the closed reports (69% to 100%) were either not classified, or their titles, labels, and content did not match the actual bugs reported. Furthermore, around 50% of the reports that did reveal bugs were not resolved by direct code patches. These findings highlight inefficiencies in the bug reporting and resolution process.
In summary, the paper classifies performance and accuracy bug reports based on symptoms, pipeline stages affected, and GitHub resolution states to better understand the bugs from deep learning frameworks.
@@ 4.1 Selection of Data Sources
We select Stack Overflow as a primary data source for our study. It is the largest programming Q&A site for programming topics containing over 24 million questions and 35 million answers SOUsageData . Thus, Stack Overflow could be a potential source for deep learning (DL) bugs. Developers often submit their encountered problems on Stack Overflow, when building their deep learning applications. According to a recent work zhao2021state , Stack Overflow contains at least 30 topics and 80K posts related to deep learning issues, which makes it a potential source for our data. We also select the benchmark dataset of Moravati et al. defects4ml (Defects4ML in Fig. 2) containing 100 DL bugs as our second source of data.
@@ 4.2 Dataset Construction
The above filtration step resulted in 3,412 Stack Overflow posts. To construct our final dataset, we apply another filter to these posts where we leverage the tags from Humbatova et al.taxonomyRealFaults . We aimed to select representative samples from different types of deep learning bugs. To reduce the risk of miscategorizing bug types, we selected tags from Table 1 that exactly matched the leaves of the bug taxonomy proposed by Humbatova et al. defects4ml . We identified several Stack Overflow tags that partially matched the leaves of the taxonomy. However, to eliminate confusion about bug types, we only used tags that had an exact match with the taxonomy. Through this manual analysis of posts, bugs and tags, we selected the tags that best represented each bug taxonomy/sub-taxonomy from Humbatova et al. taxonomyRealFaults . Table. 2 shows our selected 3 tags for model bugs, 1 for tensor bugs, 11 for training bugs, 3 for GPU bugs, and 7 for API bugs. We look for these tags in the above Stack Overflow posts and collect 113 Model bugs, 193 Tensor Bugs, 95 Training Bugs, 101 GPU Bugs, and 24 API Bugs. We also discover 42 bugs belonging to multiple categories (e.g., 5 Model & Tensor Bugs, 5 Tensor & GPU Bugs, 3 Training & API Bugs). Thus, the final dataset contains a total of 668 bugs (568 from our dataset + 100 from Defects4ML defects4ml ) and captures a balanced representation from different types of DL bugs, as shown in Table 3 (Step 3, Fig. 2).
@@ 4.4 Qualitative Analysis
Before performing any qualitative analysis, we check if each selected post accurately represents the type of bug it is assigned to. To validate the bug categorization, we manually analyze each selected Stack Overflow post and review the issue description, stack trace, observed behaviour, expected behaviour, and accepted answer. This ensures that the tag-based selection of posts from Stack Overflow does not affect our results’ validity. We perform this manual verification since the Stack Overflow posts might lack an independently verified category label. In contrast, the GitHub issues in the benchmark dataset by Morovati et al. defects4ml already contain a validated category label.
Once we confirm the category of each post or issue, we follow a two-step approach to reproduce the deep learning bugs from Stack Overflow posts and Github issues. First, we gather complementary information about each bug, such as the dataset, code snippet, library versions, and the framework used. Second, we attempt to reproduce the bug using the code snippet and supporting data (e.g., dataset information, environment configurations, hyperparameters, and training logs). The first author and one independent collaborator conducted the bug reproduction process. We first reproduced 10 bugs from our dataset and achieved a Cohen Kappa of 54.5%. Then, we had two meetings to identify the main reasons for our disagreements and resolved them. In the next round, we reproduced 10 more bugs and achieved a Cohen Kappa of 89.1%, which is considered an almost perfect agreement mchugh2012interrater . After achieving an almost perfect agreement, the first author reproduced the remaining bugs (i.e., 128) while the independent collaborator checked the reproduction, achieving an average Cohen Kappa of 85.4%. If the first author fails to reproduce the bug within 60 minutes, the independent collaborator also attempts to reproduce the bug. If both of them failed to reproduce the bug, the bug was marked as reproducible. We spent 280 person-hours on the manual bug reproduction process.
4.5 Verification of Bug Reproduction
To verify the successful reproduction of bugs manifesting in error messages, we match the error messages with those found in the bug reports. On the other hand, to verify the reproduction of bugs with silent symptoms such as slow training or low accuracy, we adopt randomness in our edit actions such as input data generation, hyperparameter initialization, and neural network construction. For example, we generate random input data from the distributions or ranges suggested in the bug reports to avoid any unintended behaviours and to address the stochastic nature of these bugs. We also construct multiple neural networks of the same architecture but with different numbers of parameters or layers. Furthermore, we run the buggy code pertaining to silent bugs multiple times with different random values to ensure that the symptoms stem from the described bug rather than our added code. Following the successful reproduction of a bug, we capture various information related to each bug, such as the deep learning architecture involved, edit actions used to reproduce the bug, time taken to reproduce the bug, type of bug reproduced, and the type of information present in the bug report. All these data gathered during bug reproduction helped us identify key edit actions and information components to reproduce specific types of deep learning bugs (Step 4, Fig. 2).